OpenStudy (anonymous):

Two dice are rolled. Find the probability that an even number is rolled on one of the dice and the number 3 is rolled on the other

OpenStudy (anonymous):

probability of a even number =3/6 probability of get a 3 =1/6 probability of both events =(1/6)(3/6) they ar independent events

OpenStudy (anonymous):

Okay THanks